OpenStudy (anonymous):

(square root of 3)sine(theta) + cosine(theta) = 1

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \sqrt{3} \times sin(x)+cos(x)=1 }\) LIKE THIS ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \sqrt{3} \times \sqrt{sin^2(x)} ~+~\sqrt{1-sin^2(x)}=0 }\) \(\Huge\color{blue}{ \sf \sqrt{3sin^2(x)} ~+~\sqrt{1-sin^2(x)}=0 }\) \(\Huge\color{blue}{ \sf \sqrt{3sin^2(x)} =-\sqrt{1-sin^2(x)} }\) \(\Huge\color{blue}{ \sf 3sin^2(x)=-(~~1-sin^2(x) ~~)}\)

OpenStudy (solomonzelman):

can you solve it now?

OpenStudy (anonymous):

oh yes, i got it now, thanks

OpenStudy (solomonzelman):

oh, no minus on the right side

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \sqrt{3sin^2(x)} =\sqrt{1-sin^2(x)} }\)

OpenStudy (solomonzelman):

(b/c I squared both sides and (-3)^2= POSITIVE 9 )

OpenStudy (solomonzelman):

and no roots, sorry. If you don't get what I meant tell me.

OpenStudy (anonymous):

im a little lost, where does the positive 9 come from?

OpenStudy (solomonzelman):

\(\Huge\color{olive}{ \it 3Sin^2(x)=1-Sin^2(x) }\)

OpenStudy (anonymous):

okay i think i can take it from here

OpenStudy (solomonzelman):

Tell me what you get as your final answer, to make sure you solve it correctly. I would prefer you to post the steps here (you can ask for any latex needed, and use equation editor, or just draw it using a drawing tool below ) but if you don't want to post your work here, it's totally fine ;)

OpenStudy (anonymous):

okay will do

OpenStudy (anonymous):

\[3\sin ^{2}(x)-1+\sin ^{2}(x)=0\]

OpenStudy (anonymous):

\[4\sin ^{2}(x)-1=0\]

OpenStudy (solomonzelman):

yeah. THEN.... add 1 to both sides divide both sides by 4 square-root both sides and take "inverse-sine" of what you get on the right side.